OpenStudy (anonymous):

How do you find k in this expression -4x^2+24x+20 can be written as a(x+h)^2+k

OpenStudy (amistre64):

you have to rewrite the left to math the form

OpenStudy (amistre64):

match the form that is :)

OpenStudy (anonymous):

could you show me how to do this problem please?

OpenStudy (amistre64):

-4x^2 +24x +20 ;factor out a -4, a = -4 -4(x^2 -6x -5) ; now complete the square for the innards to set the problem

OpenStudy (amistre64):

-4((x^2 -6x +9) -9 -5) ; combine like terms and compress the square -4((x-3)^2-14) ; distribute the -4 back into it now -4(x-3)^2 -4(-14) -4(x-3)^2 +56

OpenStudy (anonymous):

so k is 56??

OpenStudy (amistre64):

is it in the right place? if so, dbl chk my work to make sure I dint mess up :)

OpenStudy (anonymous):

what is in the right place? I really need a step by step on how to do this type of problem. I would really, really appreciate if you could help me with this.

OpenStudy (amistre64):

I just gave you a step by step the best way i know how. Other than trying to teach you how to do math all over again, this is the best I can give you for free...

OpenStudy (amistre64):

the hardest part is know how to "complete a square" ; which sounds really complex and all, but it really geometry

OpenStudy (anonymous):

want easy way to find k?

OpenStudy (anonymous):

without completing any square

OpenStudy (anonymous):

find the second coordinate of the vertex using \[-\frac{b}{2a}\] as x that will be k

OpenStudy (anonymous):

just thought i would mention it

OpenStudy (amistre64):

-4(x^2 -6x -5) ; now complete the square for the innards to set the problem focus on the innards: (x^2 -6x -5) we want to transform this into something we can use if we had an x^2 ; and 6 more of them we can align it up as: \begin{array}c x^2&x&x&x\\x\\x\\x\end{array} now to complete the square, how many spots do we need to fill in? \begin{array}c x^2&x&x&x\\x&1&2&3\\x&4&5&6\\x&7&8&9\end{array} add 9 to complete the square. x^2 +6x +9 is a complete square. BUT, we dont want to simply add 9 to the expression becasue that trows the whole thing out of whack. If we are going to add 9; we also have to subtract 9 (x^2 +6x +9) -9 -5 will keep us on the right track

OpenStudy (amistre64):

I mixed -6 and +6, but the same thing applies :)

OpenStudy (anonymous):

;yikes not to disagree with amistre but all you have to do is write \[-4(x^2-6x)+20\] \[-4(x-3)^2+k\] and you find k by replacing x in the original expression by 3

OpenStudy (amistre64):

lol .... thats not a disagreeance, its just a simplification of the process :)

OpenStudy (anonymous):

and if the 3 is not obvious use \[-\frac{b}{2a}\]

OpenStudy (amistre64):

abstract algebra is great and all; but its called abstract for a reason ;)

OpenStudy (amistre64):

what mathmatical procedure would you recommend for simply stating: (x^2 -6x) is rewritten as (x-3)^2 ? that doesnt imply alot of arm waving :)